OpenStudy (anonymous):

Prove trig identity: 2 / (sqrt3 cos x + sin x) = sec((pi/6)-x)

OpenStudy (anonymous):



OpenStudy (anonymous):

write sec in cos then use cos(a-b)=cosacosb+sinasinb simplify and u will get the solution guaranteed

OpenStudy (anonymous):

@ikram002p hey, how are you!? Could you help on my question?

OpenStudy (ikram002p):

im fine :) , wbu lol :P

OpenStudy (anonymous):

I'm good lol haven't been on here in awhile

OpenStudy (ikram002p):

yeah me too

OpenStudy (anonymous):

:) do you think you could help? I just have a few questions. I really appreciated your help last time

OpenStudy (ikram002p):

im kid off had headach lol , cant think :O sry my prain is freezing :) good luck

OpenStudy (anonymous):

I really need the help! Do you know anybody that could?

OpenStudy (anonymous):

@Mashy hey could you help me filling in the table?

OpenStudy (anonymous):

\[\sqrt{3 cosx + \sin x} \ \ \ \ ?\]

OpenStudy (anonymous):

no just sqrt3

OpenStudy (anonymous):

oh wait \[\frac{2}{\sqrt{3cosx + sinx}}\]

OpenStudy (anonymous):

plz write using equation editor :P

OpenStudy (anonymous):

Ok hold on :)

OpenStudy (anonymous):

\[2\div (\sqrt{3}\cos x +\sin x) = \sec((\Pi/6)-x)\]

OpenStudy (anonymous):

oh.. ok ok there is a root 3.. that should give you a hint ;-)

OpenStudy (anonymous):

I really need to see the steps lol I'm not really good at this :)

OpenStudy (anonymous):

I've been trying to get help for an hour lol

OpenStudy (anonymous):

do you know the identity cos(A)cos(B) + sin(A)sin(B) = ?

OpenStudy (anonymous):

= cos(A-B)

OpenStudy (anonymous):

yea.. so now.. look at the denominator.. can u convert it into that form?

OpenStudy (anonymous):

probably but I couldn't tell you how. I'm going to school in 30 min so I would really appreciate an answer as long as you explain why so I can understand. Should I solve for calculations before reasons?

OpenStudy (raden):

for me, for the bcos(x) + asin(x) can be converted into k cos(alpha - x) form. with k = sqrt(a^2 + b^2) and alpha = arctan(a/b)

OpenStudy (anonymous):

I have to fill in the table at the top. I don't know how to convert all of that

OpenStudy (raden):

i dont know how to do with that table, but you can find the relation of it with my steps :) see the equation of (sqrt3 cos x + sin x), here a = 1 and b = sqrt(3). so, k = sqrt(a^2+b^2) = sqrt(1^2 + (sqrt(3))^2) = sqrt(4) = 2 and to get alpha, use tan(alpha) =arctan(a/b) = arctan(1/sqrt(3)) = 30 degrees = pi/6 (note : a and b are positive then alpha in the first quadrant) back to equation above, see the left side : 2 / (sqrt3 cos x + sin x) = 2 / 2cos(pi/6 - x) = 1/cos(pi/6 - x) = sec(pi/6 - x) QED

OpenStudy (anonymous):

Thanks for this, but I don't think it helps because it doesn't fit in the table. WOuld this help you: http://openstudy.com/study#/updates/516cadb4e4b02ec89c5b336f

OpenStudy (raden):

i saw it, but they didnt use the table too :) i just prove from the left side, while they start from the right side.

OpenStudy (anonymous):

So would my three calculations be:1. sec(pi/6 - x) = 1/cos(pi/6-x) 2. = 1 / cos pi/6 cosx + sin pi/6 sinx 3. = 1 / (sqrt3 / 2cos x) + 1/2 sinx

OpenStudy (anonymous):

That's the answer to the calculation side?

OpenStudy (raden):

maybe, then just simplify by multiply by 2 4. => 2/sqrt(3)cosx+sinx but your table only incluads 3 rows, lol

OpenStudy (anonymous):

is there a step I could take away to add your 4th step?